OpenStudy (anonymous):

Solve 6(x + 1) - 4 = 3x + 2. A. 5/3 B. 0 C. 1 D. 4/9

OpenStudy (anonymous):

B

OpenStudy (anonymous):

B. 0

OpenStudy (anonymous):

6x+6-4=3x+2 6x+2=3x+2 6x-3x=0 3x=0 x=0

OpenStudy (anonymous):

so B.